Targeted Course Learning Outcomes: CLO 2.1

Question 1: Using the “Direct Mapped Cache”, assume that

  • 4-blocks, 1 word/block, direct mapped
  • Initial state is empty as shown below.

Index

V

Tag

Data

00

N

01

N

10

N

11

N

And the following addresses have been requested in the following sequences:

Request 1: Word Addresses 4, 5 and 7

Request 2: Word Addresses 0, 1 and 3

Request 3: Word Addresses 1, 2 and 3

Request 4: Word Addresses 3, 4 and 6

  1. Show the cache state Before and After each time that addresses have been requested.
  2. Calculate the Hit Ratio and Miss Ratio.

Hint: Follow the example in the next page and fill up all missing data accordingly and calculate the Hit and Miss ratios.

Request 1:

Word addr

Binary addr

Hit/miss

Cache block

4

100

Miss

00

5

101

??

??

7

??

??

??

Before

After

Index

V

Tag

Data

00

N

01

N

10

N

11

N

Index

V

Tag

Data

00

N

01

N

10

N

11

N

Request 2:

Word addr

Binary addr

Hit/miss

Cache block

0

??

??

??

1

??

??

??

3

??

??

??

Before

After

Index

V

Tag

Data

00

??

??

??

01

??

??

??

10

??

??

??

11

??

??

??

Index

V

Tag

Data

00

??

??

??

01

??

??

??

10

??

??

??

11

??

??

??

Hit Ratio :

Miss Ratio :
Answer #1

Total # of block inside cache = 4. Therefore index offset =2 (Least Significant 2 bits since block offset is 0)

Tag

(1 bit)

Index Offset

(2 bit)

Word Offset

(0 bits)

Requesr 1:

Word Address Binary Address Hit/miss Cache block
4 100 Miss 00
5 101 Miss 01
7 111 Miss 11

After

Index

Valid Tag Data

After

Index

Valid Tag Data
00 N 00 V 1 4
01 N 01 V 1 5
10 N 10 N
11 N 11 V 1 7

Request 2:

Word Address Binary Address Hit/miss Cache block
0 000 Miss 00
1 001 Miss 01
3 011 Miss 11

After

Index

Valid Tag Data

After

Index

Valid Tag Data
00 V 1 4 00 V 0 0
01 V 1 5 01 V 0 1
10 N 10 N
11 V 1 7 11 V 0 3

Request 3:

Word Address Binary Address Hit/miss Cache block
1 001 Hit 00
2 010 Miss 10
3 011 Hit 11

After

Index

Valid Tag Data

After

Index

Valid Tag Data
00 V 0 0 00 V 0 0
01 V 0 1 01 V 0 1
10 N 10 V 0 2
11 V 0 3 11 V 0 3

Request 4:

Word Address Binary Address Hit/miss Cache block
3 011 Hit 11
4 100 Miss 00
6 110 Miss 10

After

Index

Valid Tag Data

After

Index

Valid Tag Data
00 V 0 0 00 V 1 4
01 V 0 1 01 V 0 1
10 V 0 2 10 V 1 6
11 V 0 3 11 V 0 3

Total # of reference = 12

Total # of miss reference = 9

Total # of hit reference = 3

Therefore miss ratio= # of hit reference/ Total # of reference = 9/12=0.75

Hit ratio = # of hit reference/ Total # of reference = 3/12 =0.25

If you have any doubts, comment down. Also please upvote, thanks.